OpenStudy (ccastudentgrade10):

FAN AND MEDAL simplify the expression (x - 5)(x - 5) = 144 PLEASE HELP ASAP

OpenStudy (unklerhaukus):

simplify the left hand side (x - 5)(x - 5) = x(x - 5) -5(x - 5) =

OpenStudy (ccastudentgrade10):

i have no idea whats next

OpenStudy (ccastudentgrade10):

@UnkleRhaukus

OpenStudy (unklerhaukus):

expand x(x - 5)

OpenStudy (ccastudentgrade10):

x^2 - 5x?

OpenStudy (anonymous):

yes now do the same with 5(x-5)

OpenStudy (ccastudentgrade10):

so x^2 - 5x*x^2 - 5x

OpenStudy (ccastudentgrade10):

@leon549

OpenStudy (unklerhaukus):

not quite, what did you got when expanding 5(x-5)

OpenStudy (ccastudentgrade10):

5x - 25?

OpenStudy (unklerhaukus):

yeah, so you have (x - 5)(x - 5) = x(x - 5) -5(x - 5) = x^2 - 5x - (5x - 25) =

OpenStudy (anonymous):

(x - 5)(x - 5) = x(x - 5) -5(x - 5) = x^2 - 5x - (5x - 25) =x^2 - 5x-5x +5

OpenStudy (anonymous):

60 leon549 Medals 0 (x - 5)(x - 5) = x(x - 5) -5(x - 5) = x^2 - 5x - (5x - 25) =x^2 - 5x-5x +25

OpenStudy (ccastudentgrade10):

im very lost and i have three minutes to finish my quiz could someone explain like the whole thing in one message?? then ill do my two other problems of the same thing myself

OpenStudy (ccastudentgrade10):

how do you simplify it past what you guys have already done? @leon549 @UnkleRhaukus

OpenStudy (unklerhaukus):

combine the x terms

OpenStudy (unklerhaukus):

then you have x^2 -10x +25 = 144

OpenStudy (ccastudentgrade10):

okay so should we subtract 25?

OpenStudy (unklerhaukus):

yeah, subtract 25 from both sides

OpenStudy (ccastudentgrade10):

so x^2 - 10x = 119

OpenStudy (unklerhaukus):

then subtract 119 from both sides

OpenStudy (ccastudentgrade10):

so x^2 - 10x - 119 = 0?

OpenStudy (unklerhaukus):

you have a quadratic (x+ a )(x+ b) = 0 a * b = -119 a+b = -10

OpenStudy (unklerhaukus):

or use quadratic formula

OpenStudy (ccastudentgrade10):

so what do i do now? sorry im so out of it i was absent for a few days

OpenStudy (ccastudentgrade10):

@UnkleRhaukus

OpenStudy (unklerhaukus):

do you want to use the quadratic formula?

OpenStudy (ccastudentgrade10):

i dont know how so whatevers easiest

OpenStudy (unklerhaukus):

For a quadratic equation \[ax^2+bx+c=0\] the quadratic formula is \[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (ccastudentgrade10):

so how do i put my equation into that?

OpenStudy (unklerhaukus):

you have \[x^2 - 10x - 119 = 0\] so \(a=1\), \(b=-10\), \(c=-199\) \[x=\frac{-(-10)\pm\sqrt{(-10)^2-4(1)(-119)}}{2(1)}\\\quad=\]

OpenStudy (unklerhaukus):

(it turns out ok, i promise)

OpenStudy (ccastudentgrade10):

so x = -7 or 17?

OpenStudy (unklerhaukus):

yes!

OpenStudy (ccastudentgrade10):

thank you so much for working with me on this!!

OpenStudy (unklerhaukus):

if we knew our 17 times tables we might have gotten (x+7)(x-17) = 0 which has the same solutions as (x+7) = 0 => x = -7 or (x-17) = 0 => x = 17

OpenStudy (misty1212):

this question is not "simplify" it is "solve for x"

OpenStudy (unklerhaukus):

x = ... , is as simple as it gets.

OpenStudy (misty1212):

\[(x-5)^2=144 \\x-5=12\iff x=14\\ x-5=-12\iff x=-7\]

OpenStudy (misty1212):

ok actually \(x-5=12\iff x=17\)

OpenStudy (unklerhaukus):

oh, shortcut.

OpenStudy (unklerhaukus):

nice.

OpenStudy (misty1212):

takes two steps only and "simplify" has nothing to do with it no, it is not a shortcut, it is how you solve a quadratic equation

OpenStudy (misty1212):

thank you though! \(\color\magenta\heartsuit\)